OpenStudy (anonymous):

Help?? In 2003, Yvonne spent $5.99 on a collectible figurine. In 2008, she sold it for $35. What was the average annual percentage yield (APY) of her investment? Round to the nearest hundredth.

OpenStudy (phi):

would you use the equation P0* (1+i)^n = P where P0 is 5.99, P is 35, and n is 5 years

OpenStudy (phi):

if so, divide both sides by P0 to get \[ (1+i)^n = \frac{P}{P_0}\] raise both sides to the 1/n power to get \[1+i = \left(\frac{P}{P_0} \right)^\frac{1}{n} \]

OpenStudy (phi):

can you do this ?

OpenStudy (anonymous):

@phi So 1+i = 0.70%?

OpenStudy (phi):

what did you get for P/P0 ?

OpenStudy (anonymous):

Oh whoops. Did it backwards. I did it again.. 1.42%?

OpenStudy (phi):

better, but I would keep a few more decimals. 1.4234 so 1+i = 1.4234 what is i ?

OpenStudy (anonymous):

i=0.4234

OpenStudy (phi):

the next step is add -1 to both sides -1+1+i = 1.4234-1

OpenStudy (phi):

What was the average annual percentage yield (APY) of her investment? Round to the nearest hundredth. looks good. now change to a percent.

OpenStudy (anonymous):

thanks!! (: really helped!

OpenStudy (phi):

the only thing I would change is keep at least one more decimal place we should start with 1+i= 1.423404 (more than enough) i= 0.423404 now multiply by 100 to get 42.3404% now round to the nearest 100th 42.34%